OpenStudy (anonymous):

differentiate sin^7(4x^2+4x+3)

OpenStudy (anonymous):

\[\sin ^{7}(4x^2+4x+3)\]

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

\[7\sin ^{6}(4x ^{2}+4x+3)\cos(4x^2+4x+3)*(8x+4)\]

OpenStudy (anonymous):

i was jst testing u guys

OpenStudy (anonymous):

ow yes my bad your right XD,

OpenStudy (anonymous):

np !!!

OpenStudy (anonymous):

i dont need solution i already knw this

OpenStudy (mimi_x3):

lols HARSHIRR~ xD

OpenStudy (anonymous):

mimi x3 u r back!!!

OpenStudy (anonymous):

help me mimi virus

OpenStudy (mimi_x3):

YUPP xD

OpenStudy (mimi_x3):

LOLs , havent learnt of differentiation of sin xD

OpenStudy (anonymous):

oh !!!

OpenStudy (anonymous):

i was getting bored happy that u arrived !!!!

OpenStudy (mimi_x3):

LMAO ! im boring manss

OpenStudy (anonymous):

LMAO again !!!

OpenStudy (mimi_x3):

yuppers xD

OpenStudy (anonymous):

you will make ur aS$ contagious

OpenStudy (mimi_x3):

lmao , im willing to , so your retricewould be infected xD

OpenStudy (mimi_x3):

wtf, retrice